# Direct-CP: Directed Collaborative Perception for Connected and Autonomous Vehicles via Proactive Attention

### v2

@@ Abstract
Collaborative perception (CP) leverages visual data from connected and autonomous vehicles (CAV) to enhance an ego vehicle’s field of view (FoV). Despite recent progress, current CP methods expand the ego vehicle’s 360-degree perceptual range almost equally, which faces two key challenges. Firstly, in areas with uneven traffic distribution, focusing on directions with little traffic offers limited benefits. Secondly, under limited communication budgets, allocating excessive bandwidth to less critical directions lowers the perception accuracy in more vital areas. To address these issues, we propose Direct-CP, a proactive and direction-aware CP system aiming at improving CP in specific directions. Our key idea is to enable an ego vehicle to proactively signal its interested directions and readjust its attention to enhance local directional CP performance. To achieve this, we first propose an RSU-aided direction masking mechanism that assists an ego vehicle in identifying vital directions. Additionally, we design a direction-aware selective attention module to wisely aggregate pertinent features based on ego vehicle’s directional priorities, communication budget, and the positional data of CAVs. Moreover, we introduce a direction-weighted detection loss (DWLoss) to capture the divergence between directional CP outcomes and the ground truth, facilitating effective model training. Extensive experiments on the V2X-Sim 2.0 dataset demonstrate that our approach achieves 19.8% higher local perception accuracy in interested directions and 2.5% higher overall perception accuracy than the state-of-the-art methods in collaborative 3D object detection tasks.
@@ I Introduction
Collaborative perception (CP)[1, 2, 3] has emerged as a promising approach to expand the perceptual range of individual vehicles by integrating visual data from multiple connected and autonomous vehicles (CAVs). To effectively monitor road traffic, each CAV is equipped with an array of LiDARs or cameras that capture environmental data from various angles. This information is subsequently synthesized into a bird’s eye view (BEV) map, offering a comprehensive representation of a vehicle’s surroundings[4]. Nonetheless, relying solely on a single BEV-aided perception system is often insufficient for overcoming blind spots resulted by road obstacles or other CAVs. To address this shortcoming, CP has been adapted to allow multiple CAVs to share their local BEV features, thereby enhancing the accuracy and comprehensiveness of BEV predictions.
@@
Currently, most existing studies [5, 6] focus on optimizing 360-degree omnidirectional CP performance, aiming to extend an ego CAV’s scope in every direction almost equally. However, this overlooks the uneven traffic density across different directions and the varying interest of an ego CAV in specific directions. For instance, as illustrated in Fig. 1, when an ego CAV is making a right turn at an intersection, it may encounter minimal traffic to its rear and left front, whereas the traffic is significantly more complex to its right front. In such scenarios, the ego CAV would benefit from a targeted perception enhancement towards its right front while maintaining basic (e.g., single-vehicle) perception for other directions. Existing methods aim to uniformly enhance perception across all directions, lacking the flexibility for an ego CAV to proactively adjust its view-level priority.
@@
In addition, the communication overhead is critical factor that must be carefully considered when designing a CP system[7, 8, 2, 9, 10, 11, 12, 13]. With constraints such as a limited communication budget and a maximum allowable delay, engaging all collaborators and fully utilizing their captured views for enhancing perception across 360 degrees can significantly burden both communication and computational resources. This is particularly severe when the number of collaborators and the frame rate (measured in frames per second, FPS) are high. Indeed, reallocating communication resources from less critical directions to enhance perception in more important areas is not only strategically advantageous but also enhances CP efficiency in terms of both communication and computation.
@@
Motivated by the above observation, we propose Direct-CP, which enables an ego CAV to proactively specify its interested directions and intelligently optimize perception performance toward these directions under the constraints of a limited communication budget. To achieve this, we plan to deploy several roadside units (RSUs) to monitor the traffic distribution around the ego CAV. These RSUs provide critical data that assists the ego vehicle in determining its interested directions. Additionally, we have developed a direction-aware attention module that inputs the ego CAV’s preferred directions, communication budget, and the positional information of other CAVs, thereby generating sparse query maps that can intelligently select the most relevant information from nearby CAVs for aggregation to enhance CP performance toward selected directions. Moreover, we define a direction-weighted detection loss (DWLoss) to measure the directional perception discrepancy between prediction and ground truth. To the best of our knowledge, this is the first work designed to optimize CP based on local directional priorities. Our contributions can be summarized as follows.
@@
  • We propose a flexible CP framework named Direct-CP, which enhances perception performance towards specific directions under a limited communication budget, tailored to the proactive interests of an ego CAV.
@@
  • We design a direction-aware selective attention module to incorporate an RSU-aided direction masking mechanism, and adaptively select relevant feature data from multi-vehicle for boosting local-directional perception. Additionally, we design a direction-weighted detection loss (DWLoss) to measure the directed perception discrepancy between the outputs and the ground truth.
@@
  • We conduct extensive experiments on collaborative 3D detection tasks and the results demonstrate that our method realizes the proactive directed CP enhancement, achieving 2.5% higher overall perception accuracy and 19.8% higher local perception accuracy in the interested directions than the state-of-the-art method.
@@ II Related Works
Who2com [14] Full Average Omnidirectional
V2VNet [8] Full Average Omnidirectional
PACP [1] Full Priority-based Omnidirectional
When2com [5] Full Agent attention Omnidirectional
V2X-ViT [15] Full Self-attention Omnidirectional
Where2com [10] Sparse Location attention Omnidirectional
Direct-CP (Ours) Sparse Proactive attention Directed
@@ II-A Collaborative perception
Collaborative perception extends vehicles’ sensing capabilities beyond individual sensor limits through intermediate-stage fusion strategies [16, 17, 18, 19, 20]. While this enables feature exchange among CAVs, increasing feature dimensions and collaborator numbers demand efficient bandwidth management. Who2com [14] employs a multi-stage handshake mechanism to compress information via matching scores, while V2VNet [8] uses graph neural networks to aggregate information from nearby CAVs. These methods, however, neglect the varying importance of individual CAVs. PACP [1] addresses this limitation with a BEV-match mechanism for CAV prioritization, but it only considers agent-level priorities while ignoring view-specific importance. Furthermore, PACP’s focus on omnidirectional perception lacks the flexibility for directional perception adjustment, which is the focus of this paper.
@@ II-B Attention-based LiDAR perception
Recent advancements in LiDAR-based CP have integrated attention mechanisms to boost performance and reduce communication overhead. When2com [5] employs scaled general attention to assess correlations among different agents, reducing transmission redundancy. V2X-ViT [15] introduces the heterogeneous multi-agent attention for fusing messages across diverse agents. However, these methods require the initial transmission of full feature maps, which consumes substantial bandwidth. More recently, Where2comm [10] advances the field by utilizing sparse feature maps with location-specific and confidence-aware attention, optimizing data exchange and processing efficiency by focusing on the most relevant features. Despite its progress, Where2comm lacks the flexibility for an ego vehicle to adjust its perceptual focus based on immediate environmental demands and may not be optimal under limited communication conditions. As outlined in Table I, our proposed Direct-CP contrasts by providing a flexible and directed perception enhancement tailored to an ego vehicle’s proactive needs under limited communication constraints. This targeted approach improves data relevance and efficiency, aligning closely with real-time needs in dynamic vehicular settings.
@@ III Methodology
Figure 2: Method overview. Ego CAV combines RSU’s DAS with its interests to create direction mask. The mask, along with initial query map, neighboring CAVs’ poses, and communication budget, are input to QC-Net. QC-Net contains: (i) Direction Control Module generating direction-prioritized query confidence maps (QCMs), and (ii) Query Clipping Layer selecting top $Q_{max}\times H\times W$ queries based on QCMs under budget constraints.
Fig. 2 illustrates our method’s architecture. RSUs deployed along the roadway monitor traffic from elevated positions, providing broader views than individual vehicles. The ego CAV periodically exchanges its status with nearby RSUs to receive direction attention scores (DAS). Based on DAS and its interests, the ego vehicle masks non-essential directions during collaborative perception. Then, guided by prioritized directions, communication budget, and neighboring CAVs’ poses, it selects optimal feature map queries to maximize directed perception performance. The following subsections detail these components.
@@ III-A RSU-aided direction masking
In this paper, we leverage RSUs to help ego CAVs identify important directions. The 360-degree space around an ego CAV is divided into $N_{dir}$ local directions. Based on ego CAV’s location and speed, the corresponding RSU projects it into its captured 2D view and calculates Direction Attention Scores (DAS) for each direction. For DAS calculation, we primarily use vehicle density as an indicator, as areas with higher vehicle density typically require more attention. The DAS from RSU is represented as $\{\mathcal{S}_{r}^{i}\}_{k=1}^{N_{dir}}=\{N_{vec}^{i}\}_{k=1}^{N_{dir}}$, where $N_{vec}^{i}$ denotes the number of detected vehicles in the $i$-th direction. When RSU is unavailable, the system can alternatively use historical CP results to estimate traffic density, ensuring system robustness. This approach can be extended to incorporate additional factors such as vehicle speeds, accident history, and road conditions. The ego CAV then combines RSU’s DAS $\{\mathcal{S}_{r}^{i}\}_{i=1}^{N_{dir}}$ with its own interest weights $\{\mathcal{I}_{e}^{i}\}_{i=1}^{N_{dir}}$ to calculate the final direction mask. These interest weights can be flexibly adjusted according to the ego vehicle’s proactive needs. In cases where the weights are uniformly assigned, the direction importance relies entirely on RSU’s DAS. The final direction mask $\{\mathcal{M}_{i}\}_{i=1}^{N_{dir}}$ is calculated as follows:
@@
Our RSU-aided direction masking mechanism offers several key advantages. First, the communication between ego vehicle and RSU only involves basic information (location, speed, and DAS), ensuring minimal bandwidth usage and real-time performance. Second, the interest weight matrix provides ego vehicles with full autonomy in direction prioritization, allowing them to override RSU suggestions when necessary. Third, the dual-threshold design ($\sigma_{1}$ and $\sigma_{2}$) enables both relative and absolute traffic complexity assessment, enhancing the system’s adaptability to various scenarios.
@@ III-B Direction-aware selective attention
Consider $N$ CAVs in total in the scenario. Assume that the direction priority, the observation sets and perception supervision of the $i$-th CAV are represented as $\mathcal{M}_{i}$, $\mathcal{X}_{i}$ and $\mathcal{Y}_{i}$, respectively. The object of our considered directed collaborative perception system is to achieve the maximized perception performance toward interested directions of all agents as a function of communication budget $B$ and the number of CAVs $N$, written as:
@@
Upon receiving a 3D point cloud, the $i$-th CAV first converts the data into a BEV map. The BEV encoder $\Phi_{bev}$ processes this map to extract features, generating the feature map $\Phi_{bev}(\mathcal{X}_{i})=\mathcal{F}_{i}\in\mathbb{R}^{H\times W\times D}$, where $H$, $W$, and $D$ represent height, width, and channel dimensions, respectively. All agents project their perceptual data into a unified global coordinate system, facilitating seamless cross-agent collaboration without the need for complex coordinate transformations. The resulting feature map is fused with each other following direction-aware selective attention (DSA). The core component of DSA is the query-control net (QC-Net) taking initial query map $\mathcal{Q}_{0}\in\mathbb{R}^{H\times W\times(N-1)}$, the embedding of nearby cooperative CAV’s pose matrices $\text{PE}(\{\mathcal{P}_{i}\}_{i=2}^{N})\in\mathbb{R}^{H\times W\times(N-1)}$, the embedding of ego CAV’s direction mask $\text{DE}(\{\mathcal{M}_{k}\}_{k=1}^{N_{dir}})\in\mathbb{R}^{H\times W\times(N-1)}$, and the communication budget as input, and generates proactive binary query maps $\{\mathcal{Q}_{k}\}_{k=2}^{N}\in\mathbb{R}^{H\times W\times(N-1)}$ (value 1 means activating transmitting data in the corresponding location of BEV feature map). The communication budget $Q_{max}\in[0,1]$ is defined as the ratio of the maximum number of activated queries to the size of the query map, which satisfies:
@@
where $Q_{max}=1$ means allowing CAVs to transmit full feature map to the ego vehicle. The QC-Net consists of a three-layer MLP. The direction control module first generates query confidence map (QCM) $\{\mathcal{C}_{k}\}_{k=2}^{N}\in\mathbb{R}^{H\times W\times(N-1)}$ for each CAV, while $\mathcal{C}_{k}^{i,j}\in[0,1]$ represents the priority of the $j$-th element of the $i$-th QCM for enhancing CP in the ego vehicle’s interested directions. Assume the direction control module is denoted with $\Phi_{dcl}(\cdot)$, QCM is calculated by:
@@
Collaborative CAVs receive these query maps and compute direction-aware sparse feature maps as $\mathcal{H}_{i}=\mathcal{Q}_{i}\odot\mathcal{F}_{i}\in\mathbb{R}^{H\times W\times D}$, where $\odot$ denotes the Hadamard product of two matrices. Subsequently, each ego vehicle fuses features from multiple agents at each spatial location:
@@
where $W_{i,j}^{DSA}\in\mathbb{R}^{H\times W}$ is DSA weights assigned to the $j$-th agent by the $i$-th agent, $\text{MAttn}(\cdot)$ represents the multi-head attention at each spatial location. The fused feature map for the ego vehicle $\mathcal{F}_{i}^{out}\in\mathbb{R}^{H\times W\times D}$ is expressed as:

### v4

# Directed-CP: Directed Collaborative Perception for Connected and Autonomous Vehicles via Proactive Attention

@@ Abstract
Collaborative perception (CP) leverages visual data from connected and autonomous vehicles (CAV) to expand an ego vehicle’s field of view (FoV). Despite recent progress, current CP methods do expand the ego vehicle’s 360-degree perceptual range almost equally, but faces two key challenges. Firstly, in areas with uneven traffic distribution, focusing on directions with little traffic offers limited benefits. Secondly, under limited communication budgets, allocating excessive bandwidth to less critical directions lowers the perception accuracy in more vital areas. To address these issues, we propose Directed-CP, a proactive and direction-aware CP system aiming at improving CP in specific directions. Our key idea is to enable an ego vehicle to proactively signal its interested directions and readjust its attention to enhance local directional CP performance. To achieve this, we first propose an RSU-aided direction masking mechanism that assists an ego vehicle in identifying vital directions. Additionally, we design a direction-aware selective attention module to wisely aggregate pertinent features based on ego vehicle’s directional priorities, communication budget, and the positional data of CAVs. Moreover, we introduce a direction-weighted detection loss (DWLoss) to capture the divergence between directional CP outcomes and the ground truth, facilitating effective model training. Extensive experiments on the V2X-Sim 2.0 dataset demonstrate that our approach achieves 19.8% higher local perception accuracy in interested directions and 2.5% higher overall perception accuracy than the state-of-the-art methods in collaborative 3D object detection tasks. Codes are available at https://github.com/yihangtao/Directed-CP.git.
@@ I Introduction
Collaborative perception (CP)[fang2024pacp, hu2024fullscenedomaingeneralizationmultiagent, zhang2024smartcoopervehicularcollaborativeperception] has emerged as a promising approach to expand the perceptual range of individual vehicles by integrating visual data from multiple connected and autonomous vehicles (CAVs). To effectively monitor road traffic, each CAV is equipped with an array of LiDARs or cameras that capture environmental data from various angles. This information is subsequently synthesized into a bird’s eye view (BEV) map, offering a comprehensive representation of a vehicle’s surroundings[10160968]. Nonetheless, relying solely on a single BEV-aided perception system is often insufficient for overcoming blind spots caused by road obstacles or other CAVs. To address this shortcoming, CP has been adapted to allow multiple CAVs to share their local BEV features, thereby enhancing the accuracy and coverage of BEV predictions.
@@
Currently, most existing studies [Liu_2020_CVPR, li2021learning] focus on optimizing 360-degree omnidirectional CP performance, aiming to extend an ego CAV’s scope in every direction almost equally. However, this overlooks the non-uniform traffic density across different directions and varying interests of an ego CAV in specific directions. For instance, as illustrated in Fig. 1, when an ego CAV is making a right turn at an intersection, it may encounter minimal traffic to its rear and left front, whereas the traffic is significantly more complex to its right front. In such scenarios, the ego CAV would benefit from a targeted perception enhancement towards its right front while maintaining basic (e.g., single-vehicle) perception for other directions. Existing methods aim to uniformly enhance perception across all directions, lacking the flexibility for an ego CAV to proactively adjust its view-level priority.
@@
In addition, communication overhead is critical factor that must be carefully considered when designing a CP system[9197364, 10.1007/978-3-030-58536-5_36, hu2024fullscenedomaingeneralizationmultiagent, huCollaborativePerceptionConnected2024, hu2022wherecomm, huAgentsCoDriverLargeLanguage2024, hu2024agentscomergelargelanguagemodel, fang2022age]. With constraints such as a limited communication budget and a maximum allowable delay, engaging all collaborators and fully utilizing their captured views for enhancing perception across 360 degrees can significantly burden both communication and computational resources. This is particularly severe when the number of collaborators and the frame rate (measured in frames per second, FPS) are high. Indeed, reallocating communication resources from less critical directions to enhance perception in more important areas is not only strategically advantageous but also enhances CP efficiency in terms of communication and computation.
@@
Motivated by the above observation, we propose Directed-CP, which enables an ego CAV to proactively specify its interested directions and intelligently optimize perception performance toward these directions under constraints. To achieve this, we propose to deploy several roadside units (RSUs) to monitor the traffic distribution around an ego CAV. These RSUs provide critical data that assists the ego vehicle in determining its interested directions. Additionally, we have developed a direction-aware attention module that inputs the ego CAV’s preferred directions, communication budget, and the positional information of other CAVs, thereby generating sparse query maps that can intelligently select the most relevant information from nearby CAVs for fusion to enhance CP performance toward the selected directions. Moreover, we define a direction-weighted detection loss (DWLoss) to measure the directional perception discrepancy between prediction and ground truth. To the best of our knowledge, this is the first work designed to optimize CP based on local directional priorities. Our contributions are summarized as follows.
@@
  • We propose a flexible CP framework named Directed-CP, which enhances perception performance towards specific directions under communication budget, tailored to the proactive interests of an ego CAV.
@@
  • We design a direction-aware selective attention module to incorporate an RSU-aided direction masking mechanism, and adaptively select relevant feature data from multi-vehicle to boost local-directional perception. Additionally, we design a direction-weighted detection loss (DWLoss) to measure the directed perception discrepancy between the outputs and the ground truth.
@@
  • We conduct extensive experiments on collaborative 3D detection tasks and demonstrate that our method realizes the proactive directed CP enhancement, achieving 2.5% higher overall perception accuracy and 19.8% higher local perception accuracy in the interested directions than the state-of-the-art method.
@@ II Related Works
Who2com [liu2020who2comcollaborativeperceptionlearnable] Full Average Omnidirectional
V2VNet [10.1007/978-3-030-58536-5_36] Full Average Omnidirectional
PACP [fang2024pacp] Full Priority-based Omnidirectional
When2com [Liu_2020_CVPR] Full Agent attention Omnidirectional
V2X-ViT [10.1007/978-3-031-19842-7_7] Full Self-attention Omnidirectional
Where2com [hu2022wherecomm] Sparse Location attention Omnidirectional
Directed-CP (Ours) Sparse Proactive attention Directed
@@ II-A Collaborative perception
Collaborative perception extends vehicles’ sensing capabilities beyond single-vehicle’s limits through intermediate-stage fusion strategies [9676458, 10160546, 10160367, 10161460, 10160871]. While this enables feature exchange among CAVs, increasing feature dimensions and the number of collaborator demand efficient bandwidth management. Who2com [liu2020who2comcollaborativeperceptionlearnable] employs a multi-stage handshake mechanism to compress information via matching scores, while V2VNet [10.1007/978-3-030-58536-5_36] uses graph neural networks to aggregate information from nearby CAVs. These methods, however, neglect the varying importance of individual CAVs. PACP [fang2024pacp] addresses this limitation with a BEV-match mechanism for CAV prioritization, but it only considers agent-level priorities while ignoring view-specific importance. Furthermore, PACP focuses on omnidirectional perception but lacks the flexibility for directional perception adjustment, which is the focus of this paper.
@@ II-B Attention-based LiDAR perception
Recent advancements in LiDAR-based CP have integrated attention mechanisms to boost performance and reduce communication overhead. When2com [Liu_2020_CVPR] employs scaled general attention to assess correlations among different agents, reducing transmission redundancy. V2X-ViT [10.1007/978-3-031-19842-7_7] introduces the heterogeneous multi-agent attention for fusing messages across diverse agents. However, these methods require initial transmissions of full feature maps, which consumes substantial bandwidth. More recently, Where2comm [hu2022wherecomm] advances the field by utilizing sparse feature maps with location-specific and confidence-aware attention, optimizing data exchange and processing efficiency by focusing on the most relevant features. Despite its progress, Where2comm lacks the flexibility for an ego vehicle to adjust its perceptual focus based on immediate environmental demands and may not be effective under communication constraints. As outlined in Table I, our proposed Directed-CP contrasts by providing a flexible and directed perception enhancement tailored to an ego vehicle’s proactive needs under limited communication constraints. This targeted approach improves data relevance and efficiency, aligning closely with real-time needs under dynamic vehicular settings.
@@ III Methodology
Figure 2: Problem overview. Ego CAV combines RSU’s DAS with its interests to create direction mask. The mask, along with initial query map, neighboring CAVs’ poses, and communication budget, are input to QC-Net. QC-Net contains: (i) Direction Control Module, generating direction-prioritized query confidence maps (QCMs), and (ii) Query Clipping Layer, selecting top $Q_{max}\times H\times W$ queries based on QCMs under budget constraints.
Fig. 2 illustrates our method’s architecture. RSUs deployed along the roadway monitor traffic from elevated positions, providing broader views than individual vehicles. An ego CAV periodically exchanges its status with nearby RSUs to receive direction attention scores (DAS). Based on DAS and its interests, the ego CAV masks non-essential directions during collaborative perception. Then, guided by prioritized directions, communication budget, and neighboring CAVs’ poses, it selects optimal feature map queries to maximize directed perception performance. The following subsections detail these components.
@@ III-A RSU-aided direction masking
In this paper, we leverage RSUs to help ego CAVs identify important directions. The 360-degree space around an ego CAV is divided into $N_{dir}$ local directions. Based on ego CAV’s location and speed, the corresponding RSU projects it into its captured 2D view and calculates Direction Attention Scores (DAS) for each direction. For DAS calculation, we primarily use vehicle density as an indicator, as areas with higher vehicle densities typically require more attention. The DAS from RSU is represented as $\{\mathcal{S}_{r}^{i}\}_{k=1}^{N_{dir}}=\{N_{vec}^{i}\}_{k=1}^{N_{dir}}$, where $N_{vec}^{i}$ denotes the number of detected vehicles in the $i$-th direction. When RSU is unavailable, the system can alternatively use historical CP results to estimate traffic density, ensuring system robustness. This approach can be extended to incorporate additional factors such as vehicle speeds, accident history, and road conditions. The ego CAV then combines RSU’s DAS $\{\mathcal{S}_{r}^{i}\}_{i=1}^{N_{dir}}$ with its own interest weights $\{\mathcal{I}_{e}^{i}\}_{i=1}^{N_{dir}}$ to calculate the final direction mask. These interest weights can be flexibly adjusted according to the ego CAV’s proactive needs. In cases where the weights are uniformly assigned, the direction importance relies entirely on RSU’s DAS. The final direction mask $\{\mathcal{M}_{i}\}_{i=1}^{N_{dir}}$ is calculated as follows:
@@
Our RSU-aided direction masking mechanism offers several key advantages. First, the communication between an ego CAV and an RSU only involves basic information (location, speed, and DAS), ensuring minimal bandwidth usage and real-time performance. Second, the interest weight matrix provides ego CAVs with full autonomy of direction prioritization, allowing them to overwrite RSU suggestions when necessary. Third, the dual-threshold design ($\sigma_{1}$ and $\sigma_{2}$) enables both relative and absolute traffic complexity assessment, enhancing the system’s adaptability to various scenarios.
@@ III-B Direction-aware selective attention
Consider $N$ CAVs in total in the scenario. Assume that the direction priority, the observation sets and perception supervision of the $i$-th CAV are represented as $\mathcal{M}_{i}$, $\mathcal{X}_{i}$ and $\mathcal{Y}_{i}$, respectively. Let $\xi_{\Phi}(B,N)=(\theta^{*},\mathcal{T}^{*})$ denote the optimal parameters and transmission strategy that maximize the directed perception performance. The objective of our considered directed collaborative perception system is to achieve the maximized perception performance toward interested directions of all agents as a function of communication budget $B$ and the number of CAVs $N$, written as:
@@
Upon receiving a 3D point cloud, the $i$-th CAV first converts the data into a BEV map. The BEV encoder $\Phi_{bev}$ processes this map to extract features, generating the feature map $\Phi_{bev}(\mathcal{X}_{i})=\mathcal{F}_{i}\in\mathbb{R}^{H\times W\times D}$, where $H$, $W$, and $D$ represent height, width, and channel dimensions, respectively. All agents project their perceptual data into a unified global coordinate system, facilitating seamless cross-agent collaboration without the need of complex coordinate transformations. The resulting feature map is fused with each other following direction-aware selective attention (DSA). The core component of DSA is the query-control net (QC-Net) taking initial query map $\mathcal{Q}_{0}\in\mathbb{R}^{H\times W\times(N-1)}$, the embedding of nearby cooperative CAVs’ pose matrices $\text{PE}(\{\mathcal{P}_{i}\}_{i=2}^{N})\in\mathbb{R}^{H\times W\times(N-1)}$, the embedding of the ego CAV’s direction mask $\text{DE}(\{\mathcal{M}_{k}\}_{k=1}^{N_{dir}})\in\mathbb{R}^{H\times W\times(N-1)}$, and the communication budget as input, and generates proactive binary query maps $\{\mathcal{Q}_{k}\}_{k=2}^{N}\in\mathbb{R}^{H\times W\times(N-1)}$ (value 1 means activating transmitting data in the corresponding location of BEV feature map). The communication budget $Q_{max}\in[0,1]$ is defined as the ratio of the maximum number of activated queries to the size of the query map satisfying:
@@
where $Q_{max}=1$ means allowing CAVs to transmit full feature map to the ego CAV. The QC-Net consists of a three-layer MLP. The direction control module first generates query confidence map (QCM) $\{\mathcal{C}_{k}\}_{k=2}^{N}\in\mathbb{R}^{H\times W\times(N-1)}$ for each CAV, while $\mathcal{C}_{k}^{i,j}\in[0,1]$ represents the priority of the $j$-th element of the $i$-th QCM for enhancing CP in the ego CAV’s interested directions. Assume the direction control module is denoted with $\Phi_{dcl}(\cdot)$, QCM is calculated by:
@@
Collaborative CAVs receive these query maps and compute direction-aware sparse feature maps as $\mathcal{H}_{i}=\mathcal{Q}_{i}\odot\mathcal{F}_{i}\in\mathbb{R}^{H\times W\times D}$, where $\odot$ denotes the Hadamard product of two matrices. Subsequently, each ego CAV fuses features from multiple agents at each spatial location:
@@
where $W_{i,j}^{DSA}\in\mathbb{R}^{H\times W}$ is DSA weights assigned to the $j$-th agent by the $i$-th agent, $\text{MAttn}(\cdot)$ represents the multi-head attention at each spatial location. The fused feature map for the ego CAV $\mathcal{F}_{i}^{out}\in\mathbb{R}^{H\times W\times D}$ is expressed as: